OpenStudy (anonymous):

use the elimination method to solve the given system of equations. 4x+y=3 10x+2y=4

OpenStudy (anonymous):

first by -2, then add

Parth (parthkohli):

Or - first by 2 and subtract

OpenStudy (ajprincess):

4x+y=3-(1) 10x+2y=4-(2) (1)*2=>8x+2y=6-(3) (3)-(2)=>2x=-2 Solve for x and then substitute in (1) to find y.

OpenStudy (anonymous):

is there even a solution? my choices are A.infinitely,B.(0,10) C.(0,3) D.no solution